OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3)with a slope of -6.

OpenStudy (anonymous):

Use the Point Slope Formula: y-y\[y-y _{1}=m(x-x_{1})\]

OpenStudy (anonymous):

Where your slope is m and your coordinate point is plugged into for the y1 and x1 value.

OpenStudy (anonymous):

y + 3 = -6(x + 2) y + 2 = 6(x + 3) y - 3 = -6(x - 2) y - 2 = 6(x - 3)

OpenStudy (anonymous):

y + 3 = -6(x + 2) Notice that we have a y+3 and an x+2. This is from the sign change indicated by the formula.

OpenStudy (anonymous):

]Identify the type of correlation in the graph below.